# Sphere Shaping with Dispersion-Aware Sequence Selection for Long-Haul Transmission Systems

### 21 Dec 2023 v1

@@ V-A Optimization and Transmission Performance
Figure 7: Simulation of D-E-ESS with D-EDI computed after applying dispersion successively over $30\times 80$ km, single channel, $8$ DMB subcarriers, $110$ GBaud transmission, net rate per channel $880$ Gbit/s. Electrical SNR and GMI versus $\log_{2}(N_{S_{108}})$ at the optimal power ($4.5$ dBm), where $N_{S_{108}}$ is the number of tested sequences per $108~{}4$D-symbols. $D\text{-}E\text{-}ESS_{4}^{1}$ exhibits the highest performance in terms of GMI.
@@
Figure 8: Simulation of conventional ESS, $E\text{-}ESS_{4}^{1}$, $D\text{-}E\text{-}ESS_{4}^{1}$ and $SSFM\text{-}ESS_{4}^{1}$ with $30\times 80$ km dispersion applier over $1$ channel $8$ DMB subcarriers, and conventional ESS with $1$ channel $1$ carrier, in $110$ GBaud over $30\times 80$ km SSMF link, net rate per channel $880$ Gbit/s. (a): Electrical SNR versus Power per channel in the non-linear channel. (b): GMI versus Power per channel. $D\text{-}E\text{-}ESS_{4}^{1}$ exhibits the same performance as that of an ideal SSFM-based sequence selection.
@@
Next, in Fig. 8, we compare the performance of four single-wavelength transmission schemes: conventional ESS over a single carrier (SC) serving as a baseline, conventional ESS over DMB using $8$ subcarriers (8DMB), $D\text{-}E\text{-}ESS_{4}^{1}$ (optimal case) schemes and $E\text{-}ESS_{4}^{1}$, both in 8DMB mode. The $D\text{-}E\text{-}ESS_{4}^{1}$ shows a $0.2$ dB increase in $\mathrm{SNR}_{elec}$ and $0.03$ bit/4D-symbol gain in GMI at a power of $4.5$ dBm compared to conventional ESS with 8DMB and it shows a $0.66$ dB increase in $\mathrm{SNR}_{elec}$ and $0.16$ bit/4D in GMI at the optimal power compared to conventional ESS over a single carrier. We can also see that, even though the $E\text{-}ESS_{4}^{1}$ has a slight SNR gain, it does not counterbalance the rate loss, and eventually there is no gain in GMI with respect to conventional ESS. On the other hand, the sign-dependent-based D-E-ESS scheme maintains the GMI advantage in both short-distance and long-haul distance transmissions. Additionally, we conducted simulations involving sequence selection based on the computation of the actual NLI produced by each sequence using the Split-step Fourier method (SSFM selection) to approximate an upper bound for the performance of a sequence selection scheme. We kept the transmitter structure as 4 concatenated DMs, each equipped with $1$ flipping bit. This approach is similar to the one outlined in [27]. Each candidate sequence was transmitted over a $30\times 80$ km standard single mode fiber, single-wavelength, single-carrier system operating at $13.75$ Gbaud, without the addition of ASE noise. After transmission, we compare the NLI power of candidate sequences and select the sequence that generates the lowest NLI. This process allows us to obtain the optimal performance of sequence selection with $4$ serial DMs, and $1$ flipping bit when only considering self-phase modulation effects. From Fig. 8, we observe that SSFM-ESS and D-E-ESS exhibit almost identical performance. This observation indicates that within the analyzed system, selecting sequences based on D-EDI can be a viable alternative to the more complex SSFM approach. In other words, the variations of the waveform shaped by dispersion effects primarily dictates nonlinear distortions in high-power regions. Nonlinearity, in this context, acts as a small perturbation. Consequently, we can feasibly substitute SSFM with the dispersion-aware metric, D-EDI, for sequence selection.
@@
Figure 9: Performance analysis of $D\text{-}E\text{-}ESS_{1}^{4}$ (a): Electrical SNR and GMI versus $m_{D}$ at optimal power ($4.5$ dBm) where $m_{D}$ is the maximum number of spans used by the dispersion applier during sequence selection. Gains are maintained even when we reduce the complexity by computing D-EDI over fewer spans. (b): D-EDI and EDI versus $m_{D}$. We add the values for conventional ESS by horizontal dashed lines. Same system configuration as in Fig.7. D-EDI demonstrates a negative correlation with $\mathrm{SNR_{elec}}$.
V-B Complexity Reduction and Interpretation of the Results
@@
In the above study, we aimed to improve sequence selection by calculating the EDI at the beginning of each span along the entire transmission link. This resulted in obtaining $30$ different EDI values, which were then averaged to derive the final D-EDI. Now, we delve into the impact of the computational complexity of D-EDI on the achieved performance. To reduce the computational load of D-EDI, we explore two methods. First, we consider calculating D-EDI for only the first $m_{D}=\left\{1,\cdots,29\right\}$ spans instead of the entire transmission line. It is worth reminding that, when $m_{D}=0$, D-EDI simplifies to original EDI. Second, we consider calculating D-EDI by summing up EDIs after the signal propagates through intervals of $N_{D}>1$ spans instead of computing it at the beginning of each span ($N_{D}=1$).
@@
In Fig. 9, we explore the first method. In Fig. 9 (a), we present the performance results in terms of $\mathrm{SNR_{elec}}$ and GMI. We include the values for conventional ESS as horizontal dashed lines for comparison. SNR exhibits a steady growth as $m_{D}$ increases. The peak value is achieved when $m_{D}=29$, implying that D-EDI calculations across the entire link yield the maximum benefit. After only three spans, the nonlinear gain brought by D-EDI-based sequence selection overcomes the penalty in increased rate loss, resulting in a higher GMI than the one obtained with conventional ESS. These observations underscore the fact that in dispersion unmanaged transmission systems, nonlinear penalties are equally distributed and independent across homogeneous spans as predicted by the GN [34] and EGN models [18]. Hence, to get the most out of sequence selection, it is crucial to calculate D-EDI for the entire link. Thereafter, in Fig. 9 (b), we illustrate the EDI and D-EDI values evaluated for the selected candidate sequences for different $m_{D}$ values. EDI is evaluated over $40$ symbols ($w=40$), which is an empirical value lower than the total channel memory $682M=$ 68 (corresponding to the channel memory of a $13.75$ Gbaud transmission over the entire link ($30\times 80$ km)) as explained in Sec. III; whereas D-EDI is assessed by averaging EDIs computed with a memory of $2$ symbols ($w=2$), corresponding to the memory of a $13.75$ Gbaud transmission over a single $80$ km span. This D-EDI used for performance evaluation of the selected sequences is calculated along the entire link. Notably, As $m_{D}$ increases, D-EDI continues to decrease and hence exhibits a negative correlation with $\mathrm{SNR}_{elec}$, while EDI struggles to do so, showing that D-EDI is a better estimator of the performance of the selected sequences for multi-span transmissions.
@@
Next, we explore the second reduced-complexity D-EDI computation method. Instead of computing EDI at the start of each span, we choose to compute it for fewer spans as explained in the caption of Fig. 10. The figure shows the performance variations in terms of $\mathrm{SNR_{elec}}$ and GMI versus $N_{D}$. We observe that, with increasing $N_{D}$, SNR experiences a gradual decline. Significantly, even with $N_{D}=10$ (corresponding to EDI calculations every $800$ km), the system still outperforms conventional ESS. In this case, computing D-EDI necessitates merely $5$ evaluations, compared to $30$ when $N_{D}=1$ and resulting in a substantial reduction in computational complexity for D-EDI. This finding suggests that, for digital multi-band (DMB) transmissions over long-haul links, we can still achieve gains while using a reduced computational complexity metric for sequence selection.
@@
To conclude this subsection, we highlight two key observations from the analyses in Fig. 9 and Fig. 10 for DMB transmissions over long-haul links: first, a sequence that performs well for D-EDI over fewer spans may not necessarily maintain its effectiveness over longer links. This observation underscores the complexity of sequence selection in optical communications. It is not sufficient to assess a sequence’s suitability based solely on its performance at a single point, such as the beginning of a link or elsewhere, or at a part of the total distance. Second, identifying an optimal sequence does not require exhaustive analysis at the start of every span. Although dispersion-induced interference accumulates over longer distances, causing significant power variations and nonlinear phase fluctuations, our goal is to identify sequences that remain stable over distance. Since interference typically builds up over an extended period, involving the overlap of several symbols, it is practical to evaluate sequences at intervals spanning a few spans, rather than continuously or at every span.
V-C The Impact of Block Length and Sequence Length Variations
@@
Up to now, we have used a fixed block length of $108$ for the comparison of our proposed schemes with state-of-the-art schemes in similar conditions. It is important to note that in the case of ESS, the block length has a significant impact on the performance (SNR and GMI). Decreasing the block length leads to increased nonlinear gain but also results in greater rate loss. Consequently, the optimal block length for ESS strikes a balance between maximizing nonlinear gain and minimizing rate loss. In Fig. 11, we plot the performance of ESS schemes for four different block lengths $l=\left\{60,108,200,300\right\}$. The considered schemes are conventional ESS alongside D-E-ESS with $16$ candidate sequences and D-E-ESS with $256$ candidate sequences. With the increase of cascaded DMs and flipping bits, the number of candidate sequences can rapidly increase beyond $16$ or $256$. For example, when $4$ serial DMs, each with $3$ flipping bits, are concatenated, there are $40962^{4\times 3}=$ 4096 candidate sequences. For these cases, we randomly select $16$ or $256$ sequences for selection. In Fig. 11, for each block length and each number of tested candidate sequences, we only show the best D-E-ESS performance found after testing various configurations of cascaded DMs and flipping bits. The main observations are:
  • For the conventional ESS, as the block length increases, the SNR gradually decreases. Concurrently, the shortest block length of $60$ provides a nonlinear gain that exceeds the relatively high rate loss, resulting in the maximum achievable GMI.
  • For D-E-ESS, as the block length increases, the SNR slightly decreases. It has its maximum value at a block length of $60$ and then levels off. At a block length of $60$, the rate loss incurred by ESS itself, along with the additional rate loss from flipping bits, becomes too substantial for the achievable rate. Even though D-E-ESS exhibits a larger nonlinear gain, the final GMI falls short of what ESS offers.
  • For longer block lengths, D-E-ESS consistently outperforms ESS due to its ability to maintain a significant nonlinear gain while mitigating excessive rate loss. Among these block lengths, D-E-ESS with a block length of $300$ achieves the highest GMI.
  • In particular, when the block length is equal to $300$, D-E-ESS with $256$ candidate sequences shows a $0.08$-bit/4D GMI advantage over ESS with the same block length and a $0.04$-bit/4D advantage over ESS with a block length of $60$. D-E-ESS with $16$ selections shows similar trends with slightly lower gains.
@@
Figure 11: Electrical SNR and GMI versus DM block length. We show the optimal performance for each block length by changing the number of flipping bits and the number of cascaded DMs. Same system configuration as in Fig.7. D-E-ESS demonstrates superior performance over different block lengths.
@@
Figure 12: GMI versus selected sequence length. We show the optimal performance versus different sequence lengths, for various block lengths, by changing the number of flipping bits and the number of cascaded DMs. Same system configuration as in Fig.7. D-E-ESS demonstrates superior performance over different sequence lengths.
@@
For a thorough analysis of the performance of D-E-ESS, we vary the sequence length under identical system configuration and report the performance in Fig. 12. We explore D-E-ESS configurations with three different block lengths and vary the number of cascaded DMs, resulting in multiple sequence lengths. For each sequence length, we report the optimal performance obtained by adjusting the number of flipping bits and by performing the selection among $16$ (full lines) or $256$ (dashed lines) candidate sequences. For ease of comparison, the performance of ESS with four block lengths is also depicted on the graph using horizontal dashed lines. When the block length is set to $300$, D-E-ESS shows the best performance due to lower rate loss and comparable nonlinear behavior with respect to shorter block lengths. $256$ selections tend to perform better for longer sequences because longer sequences necessitate cascading more DMs in series, hence providing more variability in the candidates for selection. With $256$ selections, D-E-ESS can more effectively explore and utilize these possibilities. D-E-ESS significantly outperforms ESS across almost the entire range of tested sequence lengths for block lengths of $200$ and $300$. D-E-ESS with optimal sequence-length range ($100$ to $200$ 4D symbols) also surpasses the performance of ESS with block lengths of $60$ and $108$. The optimal sequence length for all four block lengths falls within the range of $100$ to $200$ 4D symbols. We explain the observed behavior by the fact that overly long sequences capture more channel memory, yet they lead to coarser selections (at a fixed number of selections to limit the complexity). On the other hand, while very short sequences offer finer selections, they only capture a small fraction of the channel memory. Consequently, for the specific transmission scenario discussed in this article, the balance point between channel memory and selection precision lies within the range of $100$ to $200$ 4D symbols.
@@
Figure 13: Electrical SNR and GMI versus DM block length. We show the optimal point versus block length by changing the number of flipping bits and the number of cascaded DMs. Same system configuration as in Fig.7, however with $5$ WDM channels. D-E-ESS demonstrates superior performance in WDM transmission scenario.
V-D WDM Transmission Performance

### 05 Mar 2024 v2

@@ VI-A Optimization and Transmission Performance
Figure 6: Simulation of D-SS with D-EDI computed after applying dispersion successively over $30\times 80$ km, single channel, $8$ DMB subcarriers, $13.75$ GBaud per subcarrier transmission, net rate per channel $880$ Gbit/s. GMI versus $\log_{2}(N_{S_{108}})$ at the optimal power ($4.5$ dBm), where $\mathrm{N_{S_{108}}}$ is the number of tested sequences per $108~{}4$D-symbols ($w=2$ for all multi-span test, corresponds to the memory of a $13.75$ Gbaud transmission over a single $80$ km span). $\mathrm{D\text{-}SS_{4}^{1}}$ exhibits the best performance.
The multi-span transmission scenario was introduced in Sec. II. For the DMB scheme, we select sequences for a single digital band and then distribute them over the eight bands. Performing sequence selection jointly for all digital bands would introduce significant complexity that we avoid in this work. In Fig. 6, we evaluate the performance of our D-SS scheme for different cascaded DMs and different numbers of flipping bits. In long-distance links, accumulated chromatic dispersion becomes apparent as its induced inter-symbol interference involves hundreds of symbols. For these links, it becomes more important to cascade multiple 1D-DMs to generate extended candidate sequences over which we will compute D-EDI. By doing so, we account for the impact of the accumulated channel memory on the waveform, thus making a better selection process. In addition, for the same complexity level as a single DM, this approach helps in reducing rate loss. As we did in the previous section, the complexity of the scheme is evaluated as $\mathrm{N_{S_{108}}}$, the number of tested candidate sequences for the generation of $108$ 4D-symbols. We calculate D-EDI by summing up EDI values of the sequences at the input of each of the $30$ spans constituting the $30\times 80$ km link ($w=2$ corresponding to the memory of a $13.75$ Gbaud transmission over a single $80$ km span). We observe that the largest GMI occurs in the configuration where four 1D-DMs are arranged in series, and each 4D-DM has one flipping bit, denoted $\mathrm{D\text{-}SS_{4}^{1}}$, and corresponding to $\log_{2}(N_{S_{108}})=4$ (i.e. $16$ candidate sequences). The GMI curve for a single DM exhibits a significantly steeper slope compared to the one for four DMs, as the rate loss increases at a slower pace in the case of cascading four DMs.
@@
Figure 7: Simulation of ESS without SS, $\mathrm{E\text{-}SS_{4}^{1}}$, $\mathrm{D\text{-}SS_{4}^{1}}$ and $\mathrm{SSFM\text{-}SS_{4}^{1}}$ with $30\times 80$ km dispersion applier over $1$ channel $8$ DMB subcarriers, and ESS without SS with $1$ channel $1$ carrier, in $110$ GBaud over $30\times 80$ km SSMF link, net rate per channel $880$ Gbit/s. Electrical SNR versus Power per channel in the non-linear channel. $\mathrm{D\text{-}SS_{4}^{1}}$ exhibits the same performance as that of an ideal SSFM-based sequence selection.
Next, in Fig. 7, we compare the performance of four single-wavelength transmission schemes: ESS without SS over a single carrier (SC) serving as a baseline, ESS without SS over DMB using $8$ subcarriers (8DMB), $\mathrm{D\text{-}SS_{4}^{1}}$ (optimal case) and $\mathrm{E\text{-}SS_{4}^{1}}$ schemes, both in 8DMB mode. The $\mathrm{D\text{-}SS_{4}^{1}}$ shows a $0.03$ bits/4D-symbol gain in GMI at a power of $4.5$ dBm compared to ESS without SS and 8DMB. It also shows a $0.16$ bits/4D in GMI at the optimal power compared to ESS without SS over a single carrier. On the other hand, the $\mathrm{E\text{-}SS_{4}^{1}}$ has no or very limited gain in GMI. Hence, the sign-dependent-based D-SS scheme maintains the GMI advantage in both short-distance and long-haul distance transmission scenarios.
@@
Additionally, we conducted simulations involving sequence selection based on the computation of the actual NLI produced by each sequence using the Split-step Fourier method (SSFM sequence selection or SSFM-SS) to approximate an upper bound for the performance of a sequence selection scheme. We kept the transmitter structure as 4 concatenated DMs, each equipped with $1$ flipping bit. This approach is similar to the one outlined in [27]. Each candidate sequence was transmitted over a $30\times 80$ km standard single mode fiber, single-wavelength, single-carrier system operating at $13.75$ Gbaud, without the addition of ASE noise. After transmission, we compare the NLI power of candidate sequences and select the sequence that generates the lowest NLI. This process allows us to obtain the optimal performance of sequence selection with $4$ serial DMs, and $1$ flipping bit when only considering self-phase modulation effects. From Fig. 7, we observe that SSFM-SS and D-SS exhibit almost identical performance. This observation indicates that within the analyzed system, selecting sequences based on D-EDI can be a viable alternative to the more complex SSFM approach. In other words, the variations of the waveform shaped by dispersion effects primarily dictate nonlinear distortions in high-power regions. Nonlinearity, in this context, acts as a small perturbation. Consequently, we can feasibly substitute SSFM with the dispersion-aware metric D-EDI to perform sequence selection with reduced complexity.
@@
Figure 8: Performance analysis of $\mathrm{D\text{-}SS_{1}^{4}}$ (a): GMI versus $\mathrm{m_{D}}$ at optimal power ($4.5$ dBm) where $\mathrm{m_{D}}$ is the maximum number of spans used by the dispersion applier during sequence selection. Gains are maintained even when we reduce the complexity by computing D-EDI over fewer spans. (b): D-EDI and EDI versus $\mathrm{m_{D}}$. We add the values for ESS without SS by horizontal dashed lines. Same system configuration as in Fig.6. D-EDI demonstrates a negative correlation with $\mathrm{SNR_{elec}}$.
VI-B Complexity Reduction and Interpretation of the Results
@@
In the above study, we aimed to improve sequence selection by calculating the EDI at the beginning of each span along the entire transmission link. This resulted in obtaining $30$ different EDI values averaged to derive the final D-EDI. Now, we delve into the impact of reducing the number of EDI computations on the achieved performance. To do so, we explore two methods. First, we consider calculating D-EDI for only the first $m_{D}=\left\{1,\cdots,29\right\}$ spans instead of the entire transmission line. It is worth reminding that when $m_{D}=0$, D-EDI simplifies to the original EDI. Second, we consider limiting the number of EDI calculations to $N_{D}<30$ by computing EDI only at the beginning of specific spans described in the caption of Fig. 9, instead of computing it at the beginning of each span ($N_{D}=30$).
@@
In Fig. 8, we explore the first method. In Fig. 8 (a), we present the performance results in terms of $\mathrm{SNR_{elec}}$ and GMI. We include the values for ESS without SS as horizontal dashed lines for comparison. SNR exhibits a steady growth as $\mathrm{m_{D}}$ increases. The peak value is achieved when $m_{D}=29$, implying that D-EDI calculations across the entire link yield the maximum benefit. After only three spans, the nonlinear gain brought by D-EDI-based sequence selection overcomes the penalty in increased rate loss, resulting in a higher GMI than the one obtained with ESS without SS. In dispersion unmanaged transmission systems, as analyzed by the GN [34] and EGN models [19], nonlinear penalties are equally distributed and independent across homogeneous spans. Consequently, as $\mathrm{m_{D}}$ increases, the calculation for D-EDI encompasses an increasingly larger number of spans. This inclusion progressively mitigates the NLI across all spans. As a result, the system performance improves linearly with the increase of $\mathrm{m_{D}}$. Hence, to get the most out of sequence selection, it is crucial to calculate D-EDI for the entire link. Furthermore, it is important to note that the calculation of D-EDI only accounts for the influence of dispersion. However, in actual optical transmission scenarios, as the transmission distance increases, NLI also escalates, subsequently impacting energy fluctuations. Consequently, the accuracy of D-EDI calculations will be generally higher when computed over a specific number of initial spans compared to a computation over an equivalent number of final spans. This higher accuracy in the early stages of transmission translates into better performance improvements following sequence selection.
Thereafter, in Fig. 8 (b), we illustrate the EDI and D-EDI values evaluated for the selected candidate sequences for different $\mathrm{m_{D}}$ values. EDI is evaluated over $40$ symbols ($w=40$), which is an empirical value lower than the total channel memory $682M=$ 68 (corresponding to the channel memory of a $13.75$ Gbaud transmission over the entire link ($30\times 80$ km)) as explained in Sec. IV, whereas D-EDI is assessed by averaging EDIs computed with a memory of $2$ symbols ($w=2$), corresponding to the memory of a $13.75$ Gbaud transmission over a single $80$ km span. This D-EDI used for performance evaluation of the selected sequences is calculated along the entire link. Notably, as $\mathrm{m_{D}}$ increases, D-EDI continues to decrease and hence exhibits a negative correlation with $\mathrm{SNR_{elec}}$, while EDI struggles to do so, showing that D-EDI is a better estimator of the performance of the selected sequences for multi-span transmissions.
@@
Next, we explore the second reduced-complexity D-EDI computation method. Instead of computing EDI at the start of each span, we choose to compute it for fewer spans specified in the caption of Fig. 9. The figure shows the performance variations in terms of GMI versus the number of required EDI calculations $N_{D}$. We observe that, with decreasing $N_{D}$, GMI experiences a gradual decline. Significantly, even with $N_{D}=5$ (corresponding to EDI calculations every $800$ km), the system still outperforms ESS without SS. In this case, the computational complexity of D-EDI is significantly reduced compared to $N_{D}=30$. This finding suggests that, for digital multi-band (DMB) transmissions over long-haul links, we can still achieve gains while using a metric with a reduced computational complexity for sequence selection.
@@
To conclude this subsection, we highlight two key observations from the analyses in Fig. 8 and Fig. 9 for DMB transmissions over long-haul links: first, a sequence that performs well for D-EDI over fewer spans may not necessarily maintain its effectiveness over longer links. This observation underscores the complexity of sequence selection in optical communications. It is not sufficient to assess the suitability of a sequence based solely on its performance at a single point in the link, such as the beginning of a link or elsewhere, or at a part of the total distance, which is also proved by Fig. 3, in which the EDI for two random sequences exhibits continuous variations over the transmission distance. Second, identifying an optimal sequence does not necessarily require an exhaustive analysis at the start of every span. Although dispersion-induced interference accumulates over longer distances, causing significant power variations and nonlinear phase fluctuations, our goal is to identify sequences that has a lower D-EDI over the entire distance. Since interference typically builds up over an extended period, involving the overlap of several symbols, it is practical to evaluate sequences at intervals spanning several spans rather than continuously or for every span.
VI-C The Impact of Block Length and Sequence Length Variations
@@
Up to now, we have used a fixed block length of $108$ for the comparison of our proposed schemes with state-of-the-art schemes in similar conditions. It is important to note that in the case of ESS, the block length has a significant impact on the performance (GMI). Decreasing the block length leads to increased nonlinear gain but also results in greater rate loss. Consequently, the optimal block length for ESS strikes a balance between maximizing nonlinear gain and minimizing rate loss. In Fig. 10, we plot the performance of ESS schemes for four different block lengths $l~{}=\left\{60,108,200,300\right\}$. The considered schemes are ESS without SS alongside D-SS with $16$ candidate sequences and D-SS with $256$ candidate sequences. With the increase of cascaded DMs and flipping bits, the number of candidate sequences can rapidly increase beyond $16$ or $256$. For example, when $4$ serial DMs, each with $3$ flipping bits, are concatenated, there are $40962^{4\times 3}=$ 4096 candidate sequences. To limit and maintain a relatively fixed complexity, we randomly select $16$ or $256$ sequences for selection to examine both low and high complexity schemes. In Fig. 10, for each block length and each number of tested candidate sequences, we only show the best D-SS performance found after testing various configurations of cascaded DMs and flipping bits. The main observations are:
@@
Figure 10: GMI versus DM block length. We show the optimal performance for each block length by changing the number of flipping bits and the number of cascaded DMs. Same system configuration as in Fig.6. D-SS demonstrates superior performance over different block lengths.
  • For the ESS without SS, as the block length decreases, the rate loss gradually increases. Concurrently, the shortest block length of $60$ provides a nonlinear gain that exceeds the relatively high rate loss, resulting in the maximum achievable GMI.
  • For D-SS, at a block length of $60$, the rate loss incurred by ESS itself, along with the additional rate loss from flipping bits, becomes too substantial for the achievable rate. Even though D-SS exhibits a larger nonlinear gain, the final GMI falls short of what ESS can offer.
  • For longer block lengths, D-SS consistently outperforms ESS due to its ability to maintain a significant nonlinear gain while mitigating excessive rate loss. Among these block lengths, D-SS with a block length of $300$ achieves the highest GMI.
  • In particular, when the block length is equal to $300$, D-SS with $256$ candidate sequences shows a $0.08$ bits/4D GMI advantage over ESS with the same block length and a $0.04$-bits/4D advantage over ESS with a block length of $60$. D-SS with $16$ selections shows similar trends with slightly lower gains.
@@
Figure 11: GMI versus selected sequence length. We show the optimal performance versus different sequence lengths, for various block lengths, by changing the number of flipping bits and the number of cascaded DMs. Same system configuration as in Fig.6. D-SS demonstrates superior performance over different sequence lengths.
@@
For a thorough analysis of the performance of D-SS, we vary the sequence length under identical system configuration and report the performance in Fig. 11. We explore D-SS configurations with three different block lengths and vary the number of cascaded DMs, resulting in multiple sequence lengths. For each sequence length, we report the optimal performance obtained by adjusting the number of flipping bits and by performing the selection among $16$ (full lines) or $256$ (dashed lines) candidate sequences. For ease of comparison, the performance of ESS with four block lengths is also depicted on the graph using horizontal dashed lines. When the block length is set to $300$, D-SS shows the best performance due to lower rate loss and comparable nonlinear behavior with respect to shorter block lengths. $256$ selections tend to perform better for longer sequences because longer sequences necessitate cascading more DMs in series, hence providing more variability in the candidates for selection. With $256$ selections, D-SS can more effectively explore and utilize these possibilities. D-SS significantly outperforms ESS across almost the entire range of tested sequence lengths for block lengths of $200$ and $300$. D-SS with optimal sequence-length range ($100$ to $200$ 4D symbols) also surpasses the performance of ESS with block lengths of $60$ and $108$. The optimal sequence length for all four block lengths falls within the range of $100$ to $200$ 4D symbols. We explain the observed behavior by the fact that overly long sequences capture more channel memory, yet they lead to coarser selections (at a fixed number of selections to limit the complexity). On the other hand, while very short sequences offer finer selections, they only capture a small fraction of the channel memory. Consequently, for the specific transmission scenario discussed in this article, the balance point between channel memory and selection precision lies within the range of $100$ to $200$ 4D symbols.
@@
Figure 12: GMI versus DM block length. We show the optimal point versus block length by changing the number of flipping bits and the number of cascaded DMs. The same system configuration as in Fig.6, however with $5$ WDM channels. D-SS demonstrates superior performance in WDM transmission scenario.
VI-D WDM Transmission Performance